OpenStudy (anonymous):

1-3c=9c+7 what are the steps?

OpenStudy (anonymous):

1-3c=9c+7 +3 +3 ___________ 1= 12c+7 -7 -7 ------------- -6=12c -------- -6 -6 c= -1\2

OpenStudy (anonymous):

-3c-9c=7-1 -12c=6 c=-1/2